OpenStudy (anonymous):

suppose f(x)=4x/3+5 and g(x)=x/3+1. solve for x when f(x)=g(x) A). x=? B). and what is the value of f(x) at that solution? please show work so I understand the steps. thanks

OpenStudy (bahrom7893):

4x/3+5 = x/3+1 4x + 15 = x + 3 3x = -12 x = -4

OpenStudy (bahrom7893):

b) f(x) = f(-4) = 4*(-4)/3 + 5 = -16/3 + 5 = 5 - 16/3 = 15/3 - 16/3 = -1/3

OpenStudy (anonymous):

thank you, did we also work out the X/3+1 part? I get what you did before that

OpenStudy (bahrom7893):

no they're just asking for f(x)

OpenStudy (anonymous):

oh. you are so helpful. there is one last part for this: Solve the inequality: f(x) is less than or equal to g(x). Write the solution set in interval notation.

OpenStudy (bahrom7893):

f(x)<=g(x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please write the solution set in interval notation:

OpenStudy (bahrom7893):

f(x)=4x/3+5 and g(x)=x/3+1 4x/3+5 <= x/3 + 1 4x + 15 <= x + 3 4x -x <= 3-15 3x <= -12 x <= -4

OpenStudy (bahrom7893):

(-infinity; 4]

OpenStudy (anonymous):

gosh, you are really helpful. Thank you so much. I actually understand the way you are working the problems.

OpenStudy (anonymous):

oh, one question, why is the is it a positive 4 in the interval notation?

OpenStudy (anonymous):

and not a negative 4?

OpenStudy (bahrom7893):

wait it was negative four, a typo.